# Data-Free Class Incremental Gesture Recognition via Synthetic Feature Sampling

### v1

@@ 5.1 Effect of Shot Size
We attributed the positive relation between shot size and performance mainly to our design of employing prototypes constructed by mean and covariance. With more training data, there is a greater likelihood of uncovering the true and robust class distribution without being influenced by outliers or unrepresentative samples. Consequently, this allows for the generation of more diverse yet still within-distribution synthetic features.
@@ 5.2 Feature Augmentation
In light of the positive relationship between performance and the size of training samples, the instinctive approach is to augment or generate additional synthetic features for classifier training [50]. However, using fewer samples to construct prototypes increases the likelihood of biasing the simulated class distribution away from the true distribution. Then, the problem becomes how can we sample the correct features from a likely biased distribution. In the feature replay method previously outlined, where the classifier trained for the previous step is available, candidate features can be inputted into the previous classifier, and samples that cannot be classified correctly are rejected.
However, the purpose now shifts towards learning the newly introduced classes instead of replaying for old classes. Thus, we developed a sampling filter strategy to discard samples confused by multiple classes, and this strategy is grounded in the assumption that the ground truth feature space for each class is compact and distinguishable, which we have reason to believe based on the earlier observation. This process involves utilizing Mahalanobis Distance to measure the distance between each sampled feature and the feature distribution of other classes. Samples are rejected if the distance falls below a threshold . After that, we reconstruct the calibrated feature space from the filtered synthetic features, allowing us to sample new and more robust synthetic features, as described in Algo. 2.
We investigate the impact of our proposed feature augmentation for new classes in the 5-shot setting. As shown in Figure. 6, with our proposed augmentation approach, there is an improvement in both global accuracy and IFM consistent across all incremental sessions.

### v2

@@ 5.1 Effect of Shot Size
The positive relationship between shot size and performance can be explained by the fact that given more training data, there is a greater likelihood of uncovering the true and robust class distribution without being influenced by outliers or unrepresentative samples. Consequently, this allows for the generation of more diverse yet still within-distribution synthetic features.
@@ 5.2 Feature Augmentation
In light of the positive relationship between performance and the size of training samples, the instinctive approach is to augment or generate additional synthetic features for classifier training. However, using less data to construct prototypes increases the likelihood of biasing the simulated class distribution away from the true distribution. Then, the problem becomes how can we sample the correct features from a likely biased distribution. In the feature replay method previously outlined, where the classifier trained for the previous step is available, candidate features can be inputted into the previous classifier, and samples that cannot be classified correctly are rejected.
However, the purpose now shifts towards learning the newly introduced classes instead of replaying for old classes. Thus, we developed a sampling filter strategy to discard samples confused by multiple classes, and this strategy is grounded in the assumption that the ground truth feature space for each class is compact and distinguishable, which we have reason to believe based on the earlier observation. This process involves utilizing Mahalanobis Distance to measure the distance between each sampled feature and the feature distribution of other classes. Samples are rejected if the distance falls below a threshold . After that, we reconstruct the calibrated feature space from the filtered synthetic features, allowing us to sample new and more robust synthetic features, as described in Alg. 2.
We investigate the impact of our proposed feature augmentation for new classes in the 5-shot setting. As shown in Figure 6, with our proposed augmentation approach, there is an improvement in both global accuracy and IFM consistent across all incremental sessions.
Comparison with Other Approaches.
We compared our results with BOAT-MI under the full-shot setting and TEEN under the 5-shot setting, as shown in Tables 3 and 4. Our 5-shot results already outperform BOAT-MI, even in the full-shot setting, which led us to skip the time-consuming 5-shot experiment with BOAT-MI. We also found our performance to be comparable to TEEN, a method specifically designed for FSCIL. For the Shrec-2017 dataset, our approach significantly outperforms TEEN in both accuracy and IFM, and for Ego-Gesture3D, we achieve a much lower IFM.